# and quantum information science

### v2

@@ 4 Definition and time evolution of quantum complexity
A quantum circuit consists of gates. A gate is an elementary unitary operation that acts simultaneously on some (typically small) number of qubits (or other qudits—$d$-level quantum systems). Every $n$-qubit unitary operation can be implemented arbitrarily precisely, by selecting an appropriate combination of gates from a universal set [15]. Quantum complexity is heuristically defined as the minimum number of gates needed to perform a certain task. Examples include mapping an initial state to a final one. Experimentally implementing a gate costs a finite amount of time. Therefore, some quantum-complexity measures reflect the duration of the optimal algorithm for performing a task.
@@
Sec. 4.1 shows that one can predict the linear growth and switchback effect from simple counting arguments in a circuit model. Section 4.2 discusses the saturation of complexity in the context of quantum many-body equilibration. In Sec. 4.3, we introduce random circuits as simple models for chaotic dynamics. With increasing depth, such circuits mimic the properties of unitary transformations selected from the Haar measure. This phenomenon can be exploited to make many of the heuristic arguments about the growth of complexity rigorous for random quantum circuits. In Sec. 4.4 and Sec. 4.5, we present two recent rigorous arguments that prove linear growth of quantum complexity for an exponentially long time. In Sec. 4.6 we revisit the problem of proving linear growth of quantum circuit complexity for exponentially long time for the time-evolution of time-independent local Hamiltonians from the perspective of computational complexity theory. This perspective will show that significant progress in theoretical computer science is necessary to even prove superpolynomial lower bounds on quantum complexity in this setting.
@@ 4.1 Time evolution of complexity, according to counting arguments about quantum circuits
Consider a fast-scrambling quantum system evolving under a unitary transformation $U(t)\coloneqq e^{-iHt}$. How does this unitary’s complexity evolve in time? We answer this question with a simple counting argument [49, 50, 1, 2].
To show how, consider a system of $K\gg 1$ qubits. [^2] We use qubits for concreteness; the argument extends to higher-dimensional qudits. Suppose that the Hamiltonian $H$ is fast-scrambling: information about initially localized perturbations spreads across the system in many-body entanglement quickly (in a time logarithmic in the system size [51, 52]). Also, suppose that $H$ is 2-local, effecting just 2-body interactions. [^3] We just need the interaction to be $k$-local, for some $k\ll K$. We are taking $k=2$ for concreteness. By discretizing time, we can regard the evolution as formed from two-qubit gates. At each time step, all the qubits are paired, and each pair is transformed by the action of a gate (Fig. 2).
@@
For how long does the unitary’s complexity grow linearly? Another counting argument answers this question, as shown in [15, Sec. 4.5.4] and [53, Sec. 8]. So far, we have reasoned about digitally simulating evolution under a 2-local Hamiltonian. Consider, instead, an arbitrary, finite, universal set of 2-qubit gates. Consider the set of all depth-$n$ circuits formed from those gates. Each circuit implements some unitary transformation. The complexity of $U(t)$ stops growing linearly once the aforementioned circuits have explored most of the volume of the group $\mathrm{SU}(2^{K})$. [^5] This claim and the ones to follow are based on populating the group manifold with $\delta$ size balls, where $\delta$ is the tolerance. The following claims therefore hold when studying approximate circuit complexity $C_{\delta}$, see section 2. The exact derivation is reviewed in [2]. This exploration is achieved by a time $t\sim\mathcal{O}(2^{2K})$. At this time, the complexity reaches a maximal value $\mathcal{C}_{\text{max}}\sim\mathcal{O}(2^{2K})$. The complexity is expected to then fluctuate around its maximal value. After a time $\sim 2^{2^{K}}$, the complexity is expected to return to near-minimal value, in a Poincaré recurrence. Figure 3 depicts these expected features of the evolution of complexity.
@@
Using this precursor, we can test how the system would behave if the perturbation $W$ acted an amount $t$ of time earlier. What is the complexity of $W(t)$? If $W$ is the identity operator, then the precursor’s complexity vanishes always: gates used to implement $U(t)$ always cancel gates used to implement $U(-t)$.
Gates cancel less if $W$ is nontrivial. Consider Fig. 4, a circuit that implements the precursor $W(t)$. The gates leftward of $W$ represent the optimal construction of $U(t)$. That construction’s first gate does not commute with the perturbation $W$. Therefore, gate 1 contributes to the complexity of $W(t)$. In contrast, gate 2 commutes with the perturbation. Therefore, the gate cancels between the circuit implementations of $U(t)$ and $U(-t)$. Hence gate 2 does not contribute to the complexity of $W(t)$. Like gate 2, gate 3 commutes with $W$. However, gate 3 does not (generally) commute with gate 1, since both act on the same qubits. Therefore, gate 3 likely contributes to the complexity of $W(t)$.
@@
Let us answer question 1. At any given layer, the $s$ infected qubits interact to some extent with the $K-s$ uninfected qubits. These interactions change the number of infected qubits by an amount $\Delta s$. We expect on probabilistic grounds [^6] At a given layer, all qubits are paired. The number of newly infected qubits is proportional to the number of previously uninfected qubits, times the probability that they are paired with infected qubits. A qubit can’t be paired with itself; hence the $K-1$.
@@
Initially, the complexity grows exponentially, as $e^{t/\ell}$. Yet the complexity is small throughout this stage, due to the $e^{-t_{*}/\ell}$ and the scrambling time’s large size. Once $t$ is large enough, we say that the perturbation has scrambled. Afterward, the complexity grows linearly, at twice the rate at which the unitary’s complexity grows, because $W(t)$ contains two unitaries. Figure 5 illustrates the delay in the complexity’s linear growth—the switchback effect—encoded in the $-t_{*}$ in the second line of Eq. (11). Holographic complexity proposals reproduce the switchback effect and the initial exponential growth of complexity (see Sec. 7.2, particularly the discussion around Eq. (123)). This fact originally motivated the complexity conjectures discussed in Sec. 7. An interesting question for future research is to investigate whether more formal arguments with random circuits can be used to reproduce the switchback effect. The discussion above focused on a simple perturbation of one qubit at one instant, as well as on 2-qubit gates. These assumptions can be relaxed [49, 50, 1, 2, 25, 54].
@@ 4.2 Complexity saturation as a late-time stage of quantum many-body equilibration
The second law of thermodynamics suggests that state complexity should grow under generic dynamics. Of course, the second law concerns entropy, not complexity, and originally described classical systems. Yet complexity seems like it ought to obey the spirit of the second law [55]. In fact, there is an analogy, with a late time saturation of complexity characterizing the final stage of quantum many-body equilibration.
[…]
@@
Which systems thermalize is difficult to predict and prove. Thermalization is often attributed to chaos, but there are many definitions of quantum chaos [12, 13, 14]. According to early thinking, a system is quantum-chaotic if it has a chaotic semiclassical limit. Later studies defined chaos through energy-level statistics modeled by Wigner–Dyson distributions. Recent work has defined chaos through properties of out-of-time-ordered correlators [56] and spectral form factors [57, 58, 59, 12], reviewed below. Some authors acknowledge challenges of characterizing quantum chaos by writing “chaos” within quotation marks and taking the term to mean nonintegrability [60]. Yet debate surrounds the definition of quantum nonintegrability, too. According to a common definition, a system is integrable if it has extensively many nontrivial conserved quantities [13].777 Consider eigendecomposing an arbitrary quantum Hamiltonian: $H=\sum_{j}E_{j}\lvert j\rangle\!\langle j\rvert$. The Hamiltonian conserves every eigenprojector $\lvert j\rangle\!\langle j\rvert$. Furthermore, the number of eigenprojectors is exponentially large in the system’s size. Yet these facts do not render $H$ integrable; if they did, every $H$ would be integrable. A conserved quantity must at least not be a simple sum of eigenprojectors in order to be considered nontrivial. To compound the ambiguity, many equilibration results follow from empirical observations, rather than from proofs, and govern specific systems, rather than general ones. We will therefore write “chaotic” within quotation marks in this subsection to communicate expectations about behaviors common to many systems called chaotic or nonintegrable.
@@
    Local equilibration: Two-point correlators decay to $\sim 1/e$ of their initial values. Time-ordered correlators relax to approximately their long-time values. Small subsystems thermalize. This equilibration occurs at the dissipation time $t_{\rm D}$ [10], also called the thermalization time. The same timescale is called the collision time if a system admits a quasiparticle description. Under certain conditions (e.g., if the interactions are highly nonlocal), $t_{\rm D}$ is constant in $K$, the number of degrees of freedom. Under other conditions, $t_{\rm D}\sim K$. For example, this latter scaling characterizes a one-dimensional system with nearest-neighbor interactions, by the Lieb–Robinson bound [61].
@@
    Quantum information scrambling: Initially localized information spreads across the system through many-body entanglement. Out-of-time-ordered correlators decay from their $O(1)$ initial values [56]. This scrambling occurs around the scrambling time, $t_{*}$. $t_{*}$ can be parametrically larger than $t_{\rm D}$, for example if there are all-to-all couplings [52]. A lower bound limits how quickly many-body entanglement can form: $t_{*}\geq\frac{\hbar}{2\pi k_{\rm B}T}\,\log(K)$ [10]. This bound saturates only if $t_{\rm D}$ is constant. In the context of semiclassical, single-particle quantum chaos, an analogue of $t_{*}$ is called the Ehrenfest time [62].
@@
    Breakdown of quantum–classical correspondence and rise in predictiveness of random-matrix theory: This stage can begin around the same time as scrambling. Hence we have already seen one name for the time when this stage takes place: the Ehrenfest time $t_{\rm E}$, also called the Thouless time $t_{\rm Th}$. This time scale varies logarithmically with the system size: $t_{\rm E}\sim\log(K)$. How this stage manifests depends on the system and measures studied. An example involves transport in, e.g., electronic systems [63, 64]. Here, an initially narrow wave packet expands to cover a classically relevant length scale around $t_{\rm E}$. The quantum system’s behavior diverges from its classical analogue; hence the name Ehrenfest time alludes to Ehrenfest’s theorem, which highlights a parallel between quantum and classical systems. In another example, at this stage, the spectral form factor stops declining and begins to ramp upward again: random-matrix theory begins to accurately predict the spectral form factor, which ceases to depend on microscopic details [65, 66].
@@
    The spectral form factor changes from increasing linearly with time to remaining constant: This change happens at the Heisenberg time $t_{\rm H}$, or plateau time [65]. Denote by $\Delta$ the average inverse gap between consecutive eigenenergies. The Heisenberg time is $t_{\rm H}\coloneqq h/\Delta\sim e^{K}$ [64].
@@
As a caveat: different authors use the same terms differently when referring to stages of quantum many-body equilibration. Also, the above list is intended to be illustrative, not comprehensive; compiling all the stages [14, 67, 68, 66, 3, 69] would require another review. For example, subtleties in the spectral form factor define other stages; see Sec. 1.1 of [67], and Sec. 1.4.3 of [66], for clear synopses and for references. Operator dynamics in Krylov space and the spread of wavefunctions require their own treatment [3, 21, 22]. Additionally, if $K$ is finite, then revivals (Poincaré recurrences and complexity revivals) occur long after times $\sim e^{K}$. Such revivals temporarily undo the equilibration but are consistent with the second law for systems of finite size.

### v3

@@ 4 Definition and time evolution of quantum complexity
A quantum circuit consists of gates. A gate is an elementary unitary operation that acts simultaneously on some (typically small) number of qubits (or other qudits—$d$-level quantum systems). Every $n$-qubit unitary operation can be implemented arbitrarily precisely, by selecting an appropriate combination of gates from a universal set nielsen2010quantum. Quantum complexity is heuristically defined as the minimum number of gates needed to perform a certain task. Examples include mapping an initial state to a final one. Experimentally implementing a gate costs a finite amount of time. Therefore, some quantum-complexity measures reflect the duration of the optimal algorithm for performing a task.
@@
Sec. 4.1 shows that one can predict the linear growth and switchback effect from simple counting arguments in a circuit model. Section 4.2 discusses the saturation of complexity in the context of quantum many-body equilibration. In Sec. 4.3, we introduce random circuits as simple models for chaotic dynamics. With increasing depth, such circuits mimic the properties of unitary transformations selected from the Haar measure. This phenomenon can be exploited to make many of the heuristic arguments about the growth of complexity rigorous for random quantum circuits. In Sec. 4.4 and Sec. 4.5, we present two recent rigorous arguments that prove linear growth of quantum complexity for an exponentially long time. In Sec. 4.6 we revisit the problem of proving linear growth of quantum circuit complexity for exponentially long time for the time-evolution of time-independent local Hamiltonians from the perspective of computational complexity theory. This perspective will show that significant progress in theoretical computer science is necessary to even prove superpolynomial (in the system’s size) lower bounds on quantum complexity in this setting.
@@ 4.1 Time evolution of complexity, according to counting arguments about quantum circuits
Consider a fast-scrambling quantum system evolving under a unitary transformation $U(t)\coloneqq e^{-iHt}$. How does this unitary’s complexity evolve in time? We answer this question with a simple counting argument Susskind:2014jwa, Brown:2016wib, Susskind:2018pmk, Chapman:2021jbh.
To show how, consider a system of $K\gg 1$ qubits. [^3] We use qubits for concreteness; the argument extends to higher-dimensional qudits. Suppose that the Hamiltonian $H$ is fast-scrambling: information about initially localized perturbations spreads across the system in many-body entanglement quickly (in a time logarithmic in the system size Susskind:2011ap, Lashkari_13_Towards). Also, suppose that $H$ is 2-local, effecting just 2-body interactions. [^4] We just need the interaction to be $k$-local, for some $k\ll K$. We are taking $k=2$ for concreteness. By discretizing time, we can regard the evolution as formed from two-qubit gates. At each time step, all the qubits are paired, and each pair is transformed by the action of a gate (Fig. 2).
@@
For how long does the unitary’s complexity grow linearly? Another counting argument answers this question, as shown in [nielsen2010quantum, Sec. 4.5.4] and [susskind2018black, Sec. 8]. So far, we have reasoned about digitally simulating evolution under a 2-local Hamiltonian. Consider, instead, an arbitrary, finite, universal set of 2-qubit gates. Consider the set of all depth-$n$ circuits formed from those gates. Each circuit implements some unitary transformation. The complexity of $U(t)$ stops growing linearly once the aforementioned circuits have explored most of the volume of the group $\mathrm{SU}(2^{K})$. [^6] This claim and the ones to follow are based on populating the group manifold with $\delta$ size balls, where $\delta$ is the tolerance. The following claims therefore hold when studying approximate circuit complexity $C_{\delta}$, see section 2. The exact derivation is reviewed in Chapman:2021jbh. This exploration is achieved by a time $t\sim\mathcal{O}(2^{2K})$. At this time, the complexity reaches a maximal value $\mathcal{C}_{\text{max}}\sim\mathcal{O}(2^{2K})$. The complexity is expected to then fluctuate around its maximal value. After a time $\sim 2^{2^{K}}$, the complexity is expected to return to near-minimal value, in a Poincaré recurrence. Figure 3 depicts these expected features of the evolution of complexity.
@@
Using this precursor, we can test how the system would behave if the perturbation $W$ acted an amount $t$ of time earlier. What is the complexity of $W(t)$? If $W$ is the identity operator, then the precursor’s complexity vanishes always: gates used to implement $U(t)$ always cancel gates used to implement U(t)U(-t).
Gates cancel less if $W$ is nontrivial. Consider Fig. 4, a circuit that implements the precursor $W(t)$. The gates leftward of $W$ represent the optimal construction of $U(t)$. That construction’s first gate does not commute with the perturbation $W$. Therefore, gate 1 contributes to the complexity of $W(t)$. In contrast, gate 2 commutes with the perturbation. Therefore, the gate cancels between the circuit implementations of $U(t)$ and U(t)U(-t). Hence gate 2 does not contribute to the complexity of $W(t)$. Like gate 2, gate 3 commutes with $W$. However, gate 3 does not (generally) commute with gate 1, since both act on the same qubits. Therefore, gate 3 likely contributes to the complexity of $W(t)$.
@@
Let us answer question 1. At any given layer, the $s$ infected qubits interact to some extent with the KsK-s uninfected qubits. These interactions change the number of infected qubits by an amount $\Delta s$. We expect on probabilistic grounds [^7] At a given layer, all qubits are paired. The number of newly infected qubits is proportional to the number of previously uninfected qubits, times the probability that they are paired with infected qubits. A qubit can’t be paired with itself; hence the K1K-1.
@@
Initially, the complexity grows exponentially, as $e^{t/\ell}$. Yet the complexity is small throughout this stage, due to the $e^{-t_{*}/\ell}$ and the scrambling time’s large size. Once $t$ is large enough, we say that the perturbation has scrambled. Afterward, the complexity grows linearly, at twice the rate at which the unitary’s complexity grows, because $W(t)$ contains two unitaries. Figure 5 illustrates the delay in the complexity’s linear growth—the switchback effect—encoded in the $t_{*}$ in the second line of Eq. (11). Holographic complexity proposals reproduce the switchback effect and the initial exponential growth of complexity (see Sec. 7.2, particularly the discussion around Eq. (123)). This fact originally motivated the complexity conjectures discussed in Sec. 7. An interesting question for future research is to investigate whether more formal arguments with random circuits can be used to reproduce the switchback effect. The discussion above focused on a simple perturbation of one qubit at one instant, as well as on 2-qubit gates. These assumptions can be relaxed Susskind:2014jwa, Brown:2016wib, Susskind:2018pmk, Chapman:2021jbh, Stanford:2014jda, Chapman:2018lsv.
@@ 4.2 Complexity saturation as a late-time stage of quantum many-body equilibration
The second law of thermodynamics suggests that state complexity should grow under generic dynamics. Of course, the second law concerns entropy, not complexity, and originally described classical systems. Yet complexity seems like it ought to obey the spirit of the second law Brown:2017jil. In fact, there is an analogy, with a late time saturation of complexity characterizing the final stage of quantum many-body equilibration.
To elaborate, we first review equilibration and thermalization. We say that a system has equilibrated if it satisfies two necessary conditions: (i) Large-scale quantities, such as temperature and volume, remain constant. (ii) No net flow of anything, such as energy or particles, enters or leaves the system. We can also define the thermalization of a system that exchanges only heat with a temperature $T$ environment as follows. Suppose that the system is classical, and consider measuring its energy. Let $p(E)$ denote the probability of obtaining the outcome $E$. The system thermalizes as $p(E)\to e^{-E/(k_{\rm B}T)}/Z_{\rm class}$. Here $k_{\rm B}$ is Boltzmann’s constant, the partition function $Z_{\rm class}\coloneqq\int dE\,\mu(E)\,e^{-E/(k_{\rm B}T)}$ normalizes the distribution, and $\mu(E)$ denotes the density of states. Now, suppose that the system is quantum and evolves under a Hamiltonian $H$. The system thermalizes as its state (density operator) approaches $e^{-H/(k_{\rm B}T)}/Z_{\rm q}$, where $Z_{\rm q}\coloneqq{\rm Tr}(e^{-H/(k_{\rm B}T)})$. This canonical state is an example of a thermal state. Different thermal states arise if the system exchanges quantities $Q_{\alpha}$ other than heat with the environment; examples include particles. In the more general case, the thermal state becomes $\propto e^{-\big(H-\sum_{\alpha}\mu_{\alpha}Q_{\alpha}\big)/(k_{\rm B}T)}$; the $\mu_{\alpha}$s denote effective chemical potentials.
@@
Which systems thermalize is difficult to predict and prove. Thermalization is often attributed to chaos, but there are many definitions of quantum chaos Haake_10_Quantum, DAlessio_16_From, Gogolin_16_Equilibration. According to early thinking, a system is quantum-chaotic if it has a chaotic semiclassical limit. Later studies defined chaos through energy-level statistics modeled by Wigner–Dyson distributions. Recent work has defined chaos through properties of out-of-time-ordered correlators Swingle_18_Unscrambling and spectral form factors Dyson_62_Statistical, Leviandier_86_Fourier, Brezin_92_Spectral, Haake_10_Quantum, reviewed below. Some authors acknowledge challenges of characterizing quantum chaos by writing “chaos” within quotation marks and taking the term to mean nonintegrability Murthy_22_Non. Yet debate surrounds the definition of quantum nonintegrability, too. According to a common definition, a system is integrable if it has extensively many nontrivial conserved quantities DAlessio_16_From. [^8] Consider eigendecomposing an arbitrary quantum Hamiltonian: $H=\sum_{j}E_{j}\lvert j\rangle\!\langle j\rvert$. The Hamiltonian conserves every eigenprojector $\lvert j\rangle\!\langle j\rvert$. Furthermore, the number of eigenprojectors is exponentially large in the system’s size. Yet these facts do not render $H$ integrable; if they did, every $H$ would be integrable. A conserved quantity must at least not be a simple sum of eigenprojectors in order to be considered nontrivial. To compound the ambiguity, many equilibration results follow from empirical observations, rather than from proofs, and govern specific systems, rather than general ones. We will therefore write “chaotic” within quotation marks in this subsection to communicate expectations about behaviors common to many systems called chaotic or nonintegrable.
@@
    Local equilibration: Two-point correlators decay to $\sim 1/e$ of their initial values. Time-ordered correlators relax to approximately their long-time values. Small subsystems thermalize. This equilibration occurs at the dissipation time $t_{\rm D}$ Maldacena:2015waa, also called the thermalization time. The same timescale is called the collision time if a system admits a quasiparticle description. Under certain conditions (e.g., if the interactions are highly nonlocal), $t_{\rm D}$ is constant in $K$, the number of degrees of freedom. Under other conditions, $t_{\rm D}\sim K$. For example, this latter scaling characterizes a one-dimensional system with nearest-neighbor interactions, by the Lieb–Robinson bound Lieb_72_Finite.
@@
    Quantum information scrambling: Initially localized information spreads across the system through many-body entanglement. Out-of-time-ordered correlators decay from their $O(1)$ initial values Swingle_18_Unscrambling. This scrambling occurs around the scrambling time, $_{*}. tt_{*}$ can be parametrically larger than $t_{\rm D}$, for example if there are all-to-all couplings Lashkari_13_Towards. A lower bound limits how quickly many-body entanglement can form: $t_{*}\geq\frac{\hbar}{2\pi k_{\rm B}T}\,\log(K)$ Maldacena:2015waa. This bound saturates only if $t_{\rm D}$ is constant. In the context of semiclassical, single-particle quantum chaos, an analogue of $t_{*}$ is called the Ehrenfest time Larkin_69_Quasiclassical.
@@
    Breakdown of quantum–classical correspondence and rise in random-matrix theory’s predictiveness: This stage can begin around the same time as scrambling. Hence we have already seen one name for the time when this stage takes place: the Ehrenfest time $t_{\rm E}$, also called the Thouless time $t_{\rm Th}$. This time scale varies logarithmically with the system size: $t_{\rm E}\sim\log(K)$. How this stage manifests depends on the system and measures studied. An example involves transport in electronic systems Aleiner_96_Divergence, Schomerus_05_Quantum: an initially narrow wave packet expands to cover a classically relevant length scale around $t_{\rm E}$. The quantum system’s behavior diverges from its classical analogue; hence the name Ehrenfest time alludes to Ehrenfest’s theorem, which highlights a parallel between quantum and classical systems. In another example, at this stage, the spectral form factor stops declining and begins to ramp upward again: random-matrix theory begins to accurately predict the spectral form factor, which ceases to depend on microscopic details Kos_18_Many, Winer_24_Spectral, 81_Maslov_Semi. [^9] Certain semiclassical behaviors of some quantum systems may extend beyond the Ehrenfest time. 61_Maslov_Quasi, 81_Chirikov_Dynamical, 81_Shepelyanskii_Quasi. This extension mirrors the ambiguities, described above, surrounding chaos. All these ambiguities motivated our formulation that the “paradigmatic system will likely pass the following mileposts.”
@@
    The spectral form factor changes from increasing linearly with time to remaining constant: This change happens at the Heisenberg time $t_{\rm H}$, or plateau time Kos_18_Many. Denote by $\Delta$ the average inverse gap between consecutive eigenenergies. The Heisenberg time is $t_{\rm H}\coloneqq h/\Delta\sim e^{K}$ Schomerus_05_Quantum.
@@
As a caveat: different authors use the same terms differently when referring to stages of quantum many-body equilibration. Also, the above list is intended to be illustrative, not comprehensive; compiling all the stages Gogolin_16_Equilibration, Hunter_18_Chaos, Bhattacharjee_24_Thermalization, Winer_24_Spectral, Nandy:2024htc, 25_VallejoFabila_Timescales would require another review. For example, subtleties in the spectral form factor define other stages; see Sec. 1.1 of Hunter_18_Chaos, and Sec. 1.4.3 of Winer_24_Spectral, for clear synopses and for references. Operator dynamics in Krylov space and the spread of wavefunctions require their own treatment Nandy:2024htc, Parker:2018yvk, Balasubramanian:2022tpr. Additionally, if $K$ is finite, then revivals (Poincaré recurrences and complexity revivals) occur long after times $\sim e^{K}$. Such revivals temporarily undo the equilibration but are consistent with the second law for systems of finite size.